OpenStudy (anonymous):

Fill in the missing number. 1/4 x ? = 5

OpenStudy (unklerhaukus):

\[1/4 \times x=5\] \[x=5 \div 1/4\] \[x=5 \times 4 =20\]

OpenStudy (unklerhaukus):

checking \[{1 \over 4} \times 20 =5\]\[5=5\]